OpenStudy (anonymous):

Use the 2nd Fundamental Thereom of Calculus to find f'(x)

OpenStudy (anonymous):

\[\int\limits_{\pi}^{\pi ^{3}} (t ^{2}+3t) dt\]

OpenStudy (shubhamsrg):

what is, if you can tell me, 2nd fundamental theorem of calculus ?

OpenStudy (shubhamsrg):

thats not the general theorem, that is a special case.. the general theorem is if f(x) = integral (h(x)) from x=p(x) to q(x) then f ' (x) = h( p(x) ) * p '(x) + h( q(x) ) * q'(x) that should help..